OpenStudy (anonymous):

What is the factored form of the equation? Factor completely. 24q^7 – 42q^4r + 36q^3r^2 – 63r^3

OpenStudy (lgbasallote):

can you see any common factor between the terms?

OpenStudy (anonymous):

24 q, 42q, and 36 q? @igbasallote

OpenStudy (anonymous):

@igbasallote, is the answer = 3(8q^7 - 14q^4r + 12q^3r^2 - 21r^3) = 3(2q^4 (4q^3 - 7r)) = 3(2q^4 + 3r^2)(4q^3 - 7r)

OpenStudy (anonymous):

@CliffSedge, can you tell me if i have it right please

OpenStudy (anonymous):

Gimme a second to look it over . . .

OpenStudy (anonymous):

This is correct. 3(2q^4 + 3r^2)(4q^3 - 7r)

OpenStudy (anonymous):

@CliffSedge, Do i only need to write that or everything I had writen?

OpenStudy (anonymous):

It depends who you're giving the answer to whether they want to see your work.

OpenStudy (anonymous):

well it says What is the factored form of the equation? Factor completely. does that mean they want all that or not?

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

Just the final product is fine then.